OpenStudy (anonymous):

y = 8x − 9 y = 4x − 1 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer.

OpenStudy (anonymous):

what is a variable that u can eliminate?

OpenStudy (anonymous):

X? I think. I'm completely lost

OpenStudy (anonymous):

treat this as an equation

OpenStudy (anonymous):

You can add or subtract to eliminate a letter (x or y).

OpenStudy (anonymous):

\[\frac{ y=8x-9 } { y=4x-1 }\]

OpenStudy (anonymous):

so can you see any thing that can be eliminated by subtraction or addition?

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

9-1??? Sorry I didn't reply I got sidetracked

OpenStudy (anonymous):

do u know wit a variable is?

OpenStudy (anonymous):

@andrea0529

OpenStudy (anonymous):

I'm sorry I really need help I suck at math pls

OpenStudy (usukidoll):

For part B, we graph these two equations y=8x-9 and y= 4x-1 then the point of intersection is when there is a solution for both of these equations

OpenStudy (usukidoll):

For part A. substitution is a better choice.

OpenStudy (anonymous):

u can eliminate y

OpenStudy (anonymous):

so it becomes \[\frac{ 8x-9 }{ 4x-1 }\]

OpenStudy (anonymous):

Why would you say that I came on here for help because I care about my education not to be insulted by rude people like you.

OpenStudy (usukidoll):

and how the heck can you solve with that?

OpenStudy (usukidoll):

anyway set these two equations together 8x-9 = 4x-1 put all variables to the left and all numbers to the right 8x-4x=-1+9 now what's 8x-4x and what's -1+9 ?

OpenStudy (anonymous):

Fine I don't want your help. You didn't hack me shut up.

OpenStudy (usukidoll):

-_- ughhhh @ganeshie8

OpenStudy (usukidoll):

@Icedragon50 stop creating drama on here. Just leave if you can't help nicely to other OS users.

OpenStudy (anonymous):

Thank you @usukidoll for actually helping me

OpenStudy (anonymous):

\[8x-y=9\]\[4x-y=1\]

OpenStudy (anonymous):

looking it this way might help

OpenStudy (usukidoll):

@saseal that works too elimination or substitution can give the same results.

OpenStudy (anonymous):

-1+9=8 thanks so much

OpenStudy (usukidoll):

forgot one part though 8x-4x=-1+9 now what's 8x-4x and what's -1+9 ? what's 8x-4x ?

OpenStudy (anonymous):

yea its easier to put in matrix form that way :)

OpenStudy (usukidoll):

true, but most of the time , matrix form versions are taught way later o_O.

OpenStudy (usukidoll):

@andrea0529 what's 8x-4x ? or an easier way to put it by factoring the x x(8-4) what's 8-4 ?

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

yea, once you go matrix you dont go back!

OpenStudy (usukidoll):

:) so now we have 4x = 8 we divide 4 from both sides so what's 8/4 ?

OpenStudy (usukidoll):

8 divided by 4 @andrea0529 then we have x by itself after that step.

OpenStudy (usukidoll):